# Extended Circular Nim

### v1

@@ 2.3 Extended circular nim with six piles
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
@@
      Assume that $n_{2}>n_{5},n_{4}\leq n_{1}$. Note that at least one of $n_{4}\oplus n_{5}\leq n_{3},n_{4}\oplus n_{3}\leq n_{5}$ and $n_{5}\oplus n_{3}\leq n_{4}$ holds. We separate into some cases.
@@
        Consider the case that $n_{4}\oplus n_{5}\leq n_{3}$ and $n_{2}+n_{4}\leq n_{1}+n_{5}$. Let $n^{\prime}_{3}=n_{4}\oplus n_{5},n^{\prime}_{0}=n_{1}+n_{5}-n_{2}-n_{4}+n^{\prime}_{3}$ , then $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n^{\prime}_{3}+n_{5},n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0,n^{\prime}_{0}\geq n^{\prime}_{3},n_{2}>n_{5},n_{1}\geq n_{4}$ and $M^{\prime}=(n^{\prime}_{0},n_{1},n_{2},n^{\prime}_{3},n_{4},n_{5})\in_{\circlearrowleft}P$. In addition, since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$, we have $n_{0}>n_{1}+n_{3}+n_{5}-n_{2}-n_{4}\geq n_{1}+n_{5}+n^{\prime}_{3}-n_{2}-n_{4}=n^{\prime}_{0}$ and $n_{3}\geq n^{\prime}_{3}$. Thus, $M^{\prime}$ is an option of $M$.
@@
        Consider the case that $n_{4}\oplus n_{3}\leq n_{5}$ and $n_{0}+n_{4}\leq n_{1}+n_{3}$. Let $n^{\prime}_{5}=n_{4}\oplus n_{3},n^{\prime}_{2}=n_{1}+n_{3}-n_{0}-n_{4}+n^{\prime}_{5}$, then $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5},n_{3}\oplus n_{4}\oplus n^{\prime}_{5}=0,n_{0}\geq n_{3},n^{\prime}_{2}\geq n^{\prime}_{5},n_{1}\geq n_{4}$ and $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})\in_{\circlearrowleft}P$. In addition, since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$, we have $n_{2}>n_{1}+n_{3}+n_{5}-n_{0}-n_{4}\geq n_{1}+n^{\prime}_{5}+n_{3}-n_{0}-n_{4}=n^{\prime}_{2}$ and $n_{5}\geq n^{\prime}_{5}$. Thus, $M^{\prime}$ is an option of $M$.
@@
        Consider the case that $n_{5}\oplus n_{3}\leq n_{4}$. Let $n^{\prime}_{4}=n_{5}\oplus n_{3}$. If $n^{\prime}_{4}+n_{0}+n_{2}\leq n_{1}+n_{3}+n_{5}$, for $n^{\prime}_{1}=n^{\prime}_{4}+n_{0}+n_{2}-n_{3}-n_{5}$, we have $n^{\prime}_{1}>n^{\prime}_{4}$ since $n_{0}+n_{2}-n_{3}-n_{5}>0$. In addition, we have $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n_{0}+n_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n_{3}+n_{5},n_{0}>n_{3},n_{2}>n_{5}$. Thus, $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n^{\prime}_{4}+n_{0}+n_{2}>n_{1}+n_{3}+n_{5}$, since $n^{\prime}_{4}\leq n_{1}$, there exist $n^{\prime}_{0},n^{\prime}_{2}$ which satisfy $n_{5}\leq n^{\prime}_{2}\leq n_{2},n_{3}\leq n^{\prime}_{0}\leq n_{0}$ and $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n_{1}+n_{3}+n_{5}$. Here, $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n^{\prime}_{4}\leq n_{1}$ also holds. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
@@
${\rm ECN}(6_{\{2\}},i)(3\leq i\leq 4)$ Regarded as $2$-pile nim
${\rm ECN}(6_{\{3\}},i)(2\leq i\leq 4)$ Regarded as $3$-pile nim
@@
2.4 Extended circular nim with seven piles
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm ECN}(7_{\{1,2\}},4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1}=n_{4},n_{2}=n_{6},n_{3}+n_{5}=n_{0}+n_{2},n_{0}=\min(\{n_{i}\})\}$.
@@
Let $N=\{M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid M\not\in_{\circlearrowleft}P\}$ The terminal position is in $P$.
Next, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})=(a,a,a+b,a+c,a,a+d,a+b)(c+d=b)\in P$. The player cannot remove all tokens from $n_{2},n_{3},n_{5},n_{6}$. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6})$ be an option of $M$. Assume that $\min(\{n^{\prime}_{i}\})=n_{0}=a$. For $M^{\prime}=(a,a,a+b^{\prime},a+c^{\prime},a,a+d^{\prime},a+b)(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ because if $b=b^{\prime}$ then $c^{\prime}+d^{\prime}<b$, if $b=c^{\prime}$ then $c^{\prime}=c=b,d=d^{\prime}=0,b^{\prime}+0=b^{\prime}<b,$ and if $b^{\prime}=d^{\prime}$ then $b+0<b^{\prime}$ or $b+0=b^{\prime}=d^{\prime},c=c^{\prime}=0$. For $M^{\prime}=(a,a,a+b,a+c^{\prime},a,a+d^{\prime},a+b^{\prime})(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ in a similar way. Also, for $M^{\prime}=(a,a,a+b^{\prime},a+c,a+d,a+b^{\prime\prime})(b^{\prime}+b^{\prime\prime}<2b,b^{\prime},b^{\prime\prime}\leq b),$ we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ because if $b^{\prime}=b^{\prime\prime}$ then $c+d=b>b^{\prime},$ if $b^{\prime}=d$ then $c>0$ or $b^{\prime\prime}<b$, and if $b^{\prime\prime}=c$ then $d>0$ or $b^{\prime}<b$.
Assume that $\min(\{n^{\prime}_{i}\})<n_{0}$. If $M^{\prime}\in_{\circlearrowleft}P$, then there exists $i$ such that $n^{\prime}_{i}=n^{\prime}_{((i+1)\bmod 7)}=n^{\prime}_{((i+4)\bmod 7)}<n_{0}$, but selecting these three piles is not legal in ${\rm ECN}(7_{1,2},4)$. Therefore, $M^{\prime}\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\in N$. We assume $n_{0}=\min(\{n_{i}\})$. Then $M=(a,a+a_{1},a+a_{2},a+a_{3},a+a_{4},a+a_{5},a+a_{6})$.
Consider the case of $a_{2}\geq a_{5}$.
@@
For the case of $a_{5}<a_{2}$, we can have a similar discussion. Therefore, $M\in N$ has an option $M^{\prime}\in_{\circlearrowleft}P$.

### v2

@@ 3 Extended circular nim with six piles
      Assume that $n_{2}>n_{5},n_{4}\leq n_{1}$. Note that at least one of $n_{4}\oplus n_{5}\leq n_{3},n_{4}\oplus n_{3}\leq n_{5}$ and $n_{5}\oplus n_{3}\leq n_{4}$ holds. We separate into five cases.
@@
        Consider the case that $n_{4}\oplus n_{5}\leq n_{3}$ and $n_{2}+n_{4}\leq n_{1}+n_{5}$. Let $n^{\prime}_{3}=n_{4}\oplus n_{5},n^{\prime}_{0}=n_{1}+n_{5}-n_{2}-n_{4}+n^{\prime}_{3}$, then $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n^{\prime}_{3}+n_{5},n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0,n^{\prime}_{3}\leq n^{\prime}_{0},n_{5}<n_{2},n_{4}\leq n_{1}$ and $M^{\prime}=(n^{\prime}_{0},n_{1},n_{2},n^{\prime}_{3},n_{4},n_{5})\in_{\circlearrowleft}P$. In addition, we have $n_{3}\geq n^{\prime}_{3}$ and $n_{0}>n_{1}+n_{3}+n_{5}-n_{2}-n_{4}\geq n_{1}+n_{5}+n^{\prime}_{3}-n_{2}-n_{4}=n^{\prime}_{0}$ since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$. Thus, $M^{\prime}$ is an option of $M$.
@@
        Consider the case that $n_{4}\oplus n_{3}\leq n_{5}$ and $n_{0}+n_{4}\leq n_{1}+n_{3}$. Let $n^{\prime}_{5}=n_{4}\oplus n_{3},n^{\prime}_{2}=n_{1}+n_{3}-n_{0}-n_{4}+n^{\prime}_{5}$, then $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5},n_{3}\oplus n_{4}\oplus n^{\prime}_{5}=0,n_{3}<n_{0},n^{\prime}_{5}\leq n^{\prime}_{2},n_{4}\leq n_{1}$ and $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})\in_{\circlearrowleft}P$. In addition, we have $n_{5}\geq n^{\prime}_{5}$ and $n_{2}>n_{1}+n_{3}+n_{5}-n_{0}-n_{4}\geq n_{1}+n^{\prime}_{5}+n_{3}-n_{0}-n_{4}=n^{\prime}_{2}$ since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$. Thus, $M^{\prime}$ is an option of $M$.
@@
        Consider the case that $n_{5}\oplus n_{3}\leq n_{4}$. Let $n^{\prime}_{4}=n_{5}\oplus n_{3}$. If $n^{\prime}_{4}+n_{0}+n_{2}\leq n_{1}+n_{3}+n_{5}$, for $n^{\prime}_{1}=n^{\prime}_{4}+n_{0}+n_{2}-n_{3}-n_{5}$, we have $n^{\prime}_{1}>n^{\prime}_{4}$ since $n_{0}+n_{2}-n_{3}-n_{5}>0$. In addition, we have $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n_{0}+n_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n_{3}+n_{5},n_{3}<n_{0},n_{5}<n_{2}$. Thus, $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n^{\prime}_{4}+n_{0}+n_{2}>n_{1}+n_{3}+n_{5}$, since $n^{\prime}_{4}\leq n_{1}$, there exist $n^{\prime}_{0},n^{\prime}_{2}$ which satisfy $n_{5}\leq n^{\prime}_{2}\leq n_{2},n_{3}\leq n^{\prime}_{0}\leq n_{0}$ and $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n_{1}+n_{3}+n_{5}$. Here, $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n^{\prime}_{4}\leq n_{1}$ also holds. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
@@
${\rm ECN}(6_{\{2\}},i)(3\leq i\leq 4)$ Regarded as two-pile nim
${\rm ECN}(6_{\{3\}},i)(2\leq i\leq 4)$ Regarded as three-pile nim
@@
4 Extended circular nim with seven piles
@@
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm ECN}(7_{\{1,2\}},4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1}=n_{4}=\min(\{n_{i}\}),n_{2}=n_{6},n_{3}+n_{5}=n_{0}+n_{2}\}$.
@@
Next, consider a position $M\in_{\circlearrowleft}P$. Without loss of generality, we assume that $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})=(a,a,a+b,a+c,a,a+d,a+b)\in P,$ where $a,b,c,d$ are nonnegative integers and $c+d=b$. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6})$ be an option of $M$. Assume that $\min(\{n^{\prime}_{i}\})=n_{0}=a$. Then, $n^{\prime}_{0}=n^{\prime}_{1}=n^{\prime}_{4}=a$ holds. Note that $\{v_{2},v_{3},v_{5},v_{6}\},\{v_{2},v_{3},v_{6}\},$ and $\{v_{2},v_{5},v_{6}\}$ are not faces, but every other subset of $\{v_{2},v_{3},v_{5},v_{6}\}$ is a face in this ruleset. There are three cases.
  • (i)
    When $M^{\prime}=(a,a,a+b^{\prime},a+c^{\prime},a,a+d^{\prime},a+b)(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$, because if $M^{\prime}\in_{\circlearrowleft}P,$ one of $a+b^{\prime}=a+b,a+c^{\prime}=a+b,$ and $a+b^{\prime}=a+d^{\prime}$ holds, but it cannot happen by the following discussion:
    • (i-a)
      When $b=b^{\prime},c^{\prime}+d^{\prime}<b$ and we have $M\not\in_{\circlearrowleft}P$.
    • (i-b)
      When $c^{\prime}=b,$ since $c+d=b,$ we have $c^{\prime}=c=b$ and $d=d^{\prime}=0$. Since $b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,$ we also have $b^{\prime}<b$. Then, $M^{\prime}=(a,a,a+b^{\prime},a+b,a,a,a+b)$ and we consider if a rotated position $(a,a,a+b,a,a,a+b^{\prime},a+b)\in P$. However, since $0+b^{\prime}<b,(a,a,a+b,a,a,a+b^{\prime},a+b)\not\in P$ holds. Therefore, $M^{\prime}\not\in_{\circlearrowleft}P$.
    • (i-c)
      When $b=b^{\prime}=d^{\prime},$ since $c+d=b,$ we have $d=d^{\prime}=b$ and $c=c^{\prime}=0$. However, it does not satisfy $b^{\prime}+c^{\prime}+d^{\prime}<b+c+d$.
    • (i-d)
      When $b>b^{\prime}=d^{\prime},M=(a,a,a+b^{\prime},a+c^{\prime},a,a+b^{\prime},a+b).$ If a rotated position $(a+c^{\prime},a,a+b^{\prime},a+b,a,a,a+b^{\prime})\in P$, then $b+0=b^{\prime}$ holds, which is a contradiction. Also, if another rotated position $(a,a+b^{\prime},a+b,a,a,a+b^{\prime},a+c^{\prime})\in P$, then $b^{\prime}=0,b=c^{\prime},$ and $0+0=b$ holds, which is a contradiction. It is easy to confirm that every other rotated and flipped positions are not elements of $P$.
  • (ii)
    For $M^{\prime}=(a,a,a+b,a+c^{\prime},a,a+d^{\prime},a+b^{\prime})(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ in a similar way.
  • (iii)
    for $M^{\prime}=(a,a,a+b^{\prime},a+c,a,a+d,a+b^{\prime\prime})(b^{\prime}+b^{\prime\prime}<2b,b^{\prime},b^{\prime\prime}\leq b),$ we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ because if $b^{\prime}=b^{\prime\prime}$ then $c+d=b>b^{\prime},$ if $b^{\prime}=d$ then $c>0$ or $b^{\prime\prime}<b$, and if $b^{\prime\prime}=c$ then $d>0$ or $b^{\prime}<b$.
Assume that $\min(\{n^{\prime}_{i}\})<n_{0}$. If $M^{\prime}\in_{\circlearrowleft}P$, then there exists $i$ such that $n^{\prime}_{i}=n^{\prime}_{((i+1)\bmod 7)}=n^{\prime}_{((i+4)\bmod 7)}<n_{0}$, but selecting these three piles in one move is not legal in ${\rm ECN}(7_{\{1,2\}},4)$. Therefore, $M^{\prime}\in N$.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\in N$. Without loss of generality, we assume $n_{0}=\min(\{n_{i}\})$. Then $M=(a,a+a_{1},a+a_{2},a+a_{3},a+a_{4},a+a_{5},a+a_{6})$, where $a_{i}\geq 0$ for every $i$.
Without loss of generality, we consider the case of $a_{2}\geq a_{5}$.
@@
Therefore, $M\in N$ has an option $M^{\prime}\in_{\circlearrowleft}P$.
For ${\rm ECN}(7_{\{1,2\}},5)$, we show the following more generalized theorem.
@@
Let $m(>1)$ be an integer and assume that $2m+1$ is a prime number. Consider a position $M=(n_{0},n_{1},\ldots,n_{2m})$ in ${\rm ECN}((2m+1)_{\{1,2,\ldots,m-1\}},2m-1)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{M\mid n_{0}=0,n_{1}=n_{2}=\cdots=n_{m-1}=n_{m}+n_{m+1}=n_{m+2}=\cdots=n_{2m}\}$.
Proof.
Let $N=\{M=(n_{0},n_{1},\ldots,n_{2m})\mid M\not\in_{\circlearrowleft}P\}$. The terminal position is in $P$.
Next, consider a position $M\in_{\circlearrowleft}P$. Without loss of generality, we assume that $M=(n_{0},n_{1},\ldots,n_{2m})\in P$. From the definition of $P$, for any position $M\in P$, $(2m-1)\times\max(\{n_{i}\})=\sum{n_{i}}$ holds.
Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ be an option of $M$. If $n^{\prime}_{i}=n_{i}$ for $i\in\{1,2,\ldots,m-1\}\cup\{m+2,\ldots,2m\}$, then $\max(\{n_{i}\})=\max(\{n^{\prime}_{i}\})$, but $\sum{n^{\prime}_{i}}<\sum{n_{i}}$. Therefore, $(2m-1)\times\max(\{n^{\prime}_{i}\})>\sum{n^{\prime}_{i}}$ and $M^{\prime}\not\in_{\circlearrowleft}P$.
If $n^{\prime}_{i}<n_{i}$ for every $i\in\{1,\ldots,m-1\}\cup\{m+2,\ldots,2m\}$, then $n^{\prime}_{m}=n_{m}$ and $n^{\prime}_{m+1}=n_{m+1}$ hold.
Since $n^{\prime}_{m}+n^{\prime}_{m+1}=n_{m}+n_{m+1}=n_{1}>n^{\prime}_{1}$, we do not have $n^{\prime}_{0}=0,n^{\prime}_{1}=\cdots=n^{\prime}_{m-1}=n^{\prime}_{m}+n^{\prime}_{m+1}=n^{\prime}_{m+2}=\cdots=n^{\prime}_{2m}$.
Also, we do not have $n^{\prime}_{i}=0,n^{\prime}_{(i+1)\bmod(2m+1)}=\cdots=n^{\prime}_{(i+m-1)\bmod(2m+1)}=n^{\prime}_{(i+m)\bmod(2m+1)}+n^{\prime}_{(i+m+1)\bmod(2m+1)}=n^{\prime}_{(i+m+2)\bmod(2m+1)}=\cdots=n^{\prime}_{(i+2m)\bmod(2m+1)}$ for $i\in\{1,\ldots,m-1\}\cup\{m+2,\ldots,2m\}$, since in this case, we have $n^{\prime}_{0}\geq n^{\prime}_{m}$ and $n^{\prime}_{0}\geq n^{\prime}_{m+1}$, which contradicts $n^{\prime}_{0}=n_{0}=0$ and $n^{\prime}_{m}+n^{\prime}_{m+1}>0$.
The case $n^{\prime}_{m}=0,n^{\prime}_{m+1}=n^{\prime}_{m+2}=\cdots=n^{\prime}_{2m}+n^{\prime}_{0}=n^{\prime}_{1}=\cdots=n^{\prime}_{m-1}$ also contradicts because $n_{m+1}^{\prime}=n^{\prime}_{m}+n^{\prime}_{m+1}=n_{m}+n_{m+1}=n_{m+2}>n_{m+2}^{\prime}$. In a similar way, $n^{\prime}_{m+1}=0,n^{\prime}_{m+2}=n^{\prime}_{m+3}=\cdots=n^{\prime}_{2m}=n^{\prime}_{0}+n^{\prime}_{1}=n^{\prime}_{2}=\cdots=n^{\prime}_{m}$ contradicts, too.
Therefore, for $M\in P,$ every option $M^{\prime}\in N$.
Finally, consider a position $M=(n_{0},n_{1},\ldots,n_{2m})\not\in N$.
Note that except for $V\setminus\{v_{i},v_{(i+m)\bmod(2m+1)}\}(0\leq i\leq 2m),$ every set of five vertices can be a face of this ruleset since $2m+1$ is a prime number.
Without loss of generality, we assume that $\min(\{n_{i}\})=n_{0}$. We consider two cases:
  • (i)
    Without loss of generality, we assume that $\min(\{n_{1},n_{2},\ldots,n_{2m}\})=n_{i}(i\in\{1,2,\ldots,m-1\})$ or $\min(\{n_{1},n_{2},\ldots,n_{2m}\})=n_{m}$, and we consider the first case. Let $n^{\prime}_{0}=n_{0},n^{\prime}_{m}=0,n^{\prime}_{2m}=n_{i}-n_{0},$ and $n^{\prime}_{j}=n_{i}$ for any $j\in\{1,2,\ldots,m-1,m+1,\ldots,2m-1\}$.
    Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
  • (ii)
    For the second case, we assume that $\min(\{n_{1},n_{2},\ldots,n_{2m}\})=n_{m}$.
    • (ii-a)
      If $\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m}\})=n_{i}(i\in\{1,2,\ldots,2m-1\}\setminus\{m,m+1\})$, then let $n^{\prime}_{0}=0,n^{\prime}_{m}=n_{m},n^{\prime}_{m+1}=n_{i}-n_{m}$ and $n^{\prime}_{j}=n_{i}$ for any $j\in\{1,2,\ldots,2m+1\}\setminus\{0,m,m+1\}$. Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in P$.
    • (ii-b)
      If $\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m}\})=n_{2m}$ and $n_{0}+n_{2m}\leq\min(\{n_{1},n_{2},\ldots,n_{2m-1}\}\setminus\{n_{m}\})$, then let $n^{\prime}_{0}=n_{0},n^{\prime}_{m}=0,n^{\prime}_{2m}=n_{2m}$ and $n^{\prime}_{j}=n_{0}+n_{2m}$ for any $j\in\{1,\ldots,2m-1\}\setminus\{m\}$. Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (ii-c)
      Consider the case that $\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m}\})=n_{2m}$ and $n_{0}+n_{2m}>\min(\{n_{1},n_{2},\ldots,n_{2m-1}\}\setminus\{n_{m}\})$. Let $a=\min(\{n_{1},n_{2},\ldots,n_{2m-1}\}\setminus\{n_{m}\})$. $n_{0}\leq a,n_{2m}\leq a,$ and $n_{0}+n_{2m}>a$ hold.
      • (ii-c-1)
        If $n_{i}=a$ for $i\in\{1,2,\ldots,2m-1\}\setminus\{m,m+1\},$ then let $n^{\prime}_{0}=n_{0},n^{\prime}_{m}=0,n^{\prime}_{2m}=a-n_{0}$, and $n^{\prime}_{j}=a$ for $j\in\{1,2,\ldots,2m-1\}\setminus\{m\}.$ Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (ii-c-2)
        If $n_{m+1}=a$, then let $n^{\prime}_{0}=a-n_{2m},n^{\prime}_{m}=0,n^{\prime}_{2m}=n_{2m},$ and $n^{\prime}_{j}=a$ for any $j\in\{1,2,\ldots,2m-1\}\setminus\{m\}$. Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (ii-d)
      If $\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m}\})=n_{m+1}$ and $n_{m}+n_{m+1}\leq\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m},n_{m+1}\})$, then let $n^{\prime}_{0}=0,n^{\prime}_{m}=n^{\prime}_{m},n^{\prime}_{m+1}=n^{\prime}_{m+1},$ and $n^{\prime}_{j}=n_{m}+n_{m+1}$ for $j\in\{1,2,\ldots,2m\}\setminus\{m,m+1\}.$ Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in P$.
    • (ii-e)
      Consider the case that $\min(\{n_{1},n_{2},\ldots,n_{2m}\})=n_{m+1}$ and $n_{m}+n_{m+1}>\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m},n_{m+1}\})$. Let $a=\min(\{n_{1},n_{2},\ldots,n_{2m}\}\setminus\{n_{m},n_{m+1}\})$. $n_{m}\leq a,n_{m+1}\leq a$, and $n_{m}+n_{m+1}>a$ hold.
      • (ii-e-1)
        If $n_{i}=a$ for $i\in\{2,3,\ldots,2m\}\setminus\{m,m+1\},$ then let $n^{\prime}_{0}=0,n^{\prime}_{m}=a-n_{m+1},n^{\prime}_{m+1}=n_{m+1},$ and $n^{\prime}_{j}=a$ for $j\in\{1,2,\ldots,2m\}\setminus\{m,m+1\}$. Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in P$.
      • (ii-e-2)
        If $n_{1}=a$, then let $n^{\prime}_{0}=0,n^{\prime}_{m}=n_{m},n^{\prime}_{m+1}=a-n_{m},$ and $n^{\prime}_{j}=a$ for any $j\in\{1,2,\ldots,2m\}\setminus\{m,m+1\}.$ Then, $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},\ldots,n^{\prime}_{2m})$ is an option of $M$ and $M^{\prime}\in P$.
Corollary 2.